salt ph

1.      (25 pt.) I have a beaker with 250.0 mL of distilled water.  I dissolve 10.0 g of sodium acetate (NaOAc) and 10.0 g of potassium cyanate (KCNO).  What is the pH of the resulting solution?
Answer #1

They are both salts.  I dissociate the NaOAC and I get Na+ and OAc-

I dissociate the KCNO and I get K+ and CNO-


The Na+, K+ are spectators, so I ignore them.


The OAc- and the CNO- form a buffer.  If I convert g to moles, I have 0.125 moles of OAC- and 0.177 moles of CNO-


pH=pKa + log (acid/base) = 4.26 + log(0.125/0.177)= 4.11

pH=4.11!

source: instructor answer key
answered by: Tommy Cain